Find one-half of these fractions.

1. 1/4

2. 2/3

3. 3/4




1
Expert's answer
2021-07-18T14:56:03-0400

1) "\frac{1}{4}*\frac{1}{2}=\frac{1}{8}."

2) "\frac{2}{3}*\frac{1}{2}=\frac{1}{3}."

3) "\frac{3}{4}*\frac{1}{2}=\frac{3}{8}."